i dont understand -17-2x=6-x
Answer by ReadingBoosters(3246)

Starting with
-17 - 2x = 6 - x
Group like terms
-17 - 2x + 2x = 6 - x + 2x
-17 = 6 - x + 2x
-17 - 6 = 6 - 6 - x + 2x
-17 - 6 = -x + 2x
Rewrite as addition
-17 + -6 = -x + 2x
-23 = x
Post more questions if you need further clarity.
...............
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.